OpenStudy (anonymous):

City A and City B had two different temperatures on a particular day. On that day, four times the temperature of City A was 8° C more than 3 times the temperature of City B. The temperature of City A minus twice the temperature of City B was –3° C. What was the temperature of City A and City B on that day? City A was 5° C, and City B was 4° C. City A was 3° C, and City B was –1° C. City A was 8° C, and City B was –3° C. City A was 5° C, and City B was –5° C.

OpenStudy (anonymous):

Answer is: "City A was 5° C, and City B was 4° C." Solution: 4A=3B+8 A-2B=-3 4A-3B=8 4A-8B=-12 ------------- 5B=20 B=4 degree C A=5 degree C

OpenStudy (anonymous):

Thanks :)